# AiGet: Transforming Everyday Moments into Hidden Knowledge Discovery with AI Assistance on Smart Glasses

### v1

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713953isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Human-centered computing Empirical studies in interaction design
Figure 1. A day with AiGet, a wearable knowledge discovery assistant equipped with AR smart glasses and a ring mouse. (1) While casually walking, AiGet actively analyzes the user’s gaze patterns and identifies “interesting” environmental entities. (2) AiGet recognizes the spider lily that the user briefly glanced at as a valuable opportunity for informal learning, considering her interest in plants in the User Profile and lack of knowledge about the flower. (3) Using contextual data, the system generates multiple knowledge candidates and selects the most valuable one about the spider lily to the user. (4) AiGet delivers multimodal feedback, including audio, text keywords with emojis, and a image with a bounding box for the flower, to help the user locate the flower and gain information with minimal interference to primary tasks. (5) Surprised by the knowledge delivery about the overlooked flower, the user uses the ring mouse to ask follow-up questions. (6) Later, AiGet continuously helps her uncover design details behind the campus architecture and compare two unfamiliar snacks, aligning with her interest in healthy diets, etc. (7) When the user meets a friend who owns a cat, she shares what she learned from AiGet—that spider lilies are toxic to pets—enriching their social interaction.
@@ 1. Introduction
Our daily surroundings are filled with rich knowledge, from building designs to local flora and fauna, offering countless opportunities for discovery and informal learning from everyday moments (Kaplan, 1989; Bekerman et al., 2006; Ardoin and Heimlich, 2021). Such informal learning not only allows us to gather useful knowledge about our surroundings but also fosters critical thinking and diverse perspectives, helping us adapt to new situations and inspire innovation (Marsick and Watkins, 2001; Cross, 2007; Eraut *, 2004). Yet, modern adult life often disconnects us from these rich opportunities, making us “blind” to such discoveries. Pressed by daily demands, adults focus on primary tasks with higher priority (e.g., rushing for work and studies), bypassing opportunities for curiosity and exploration. We also tend to prioritize immediate rewards, avoiding the perceived effort of learning (Liquin and Gopnik, 2022). Moreover, the rise of smartphones and social media, which offer effortless knowledge, has further diverted our attention from the physical world. As a result, we may overlook rare migratory birds in our backyard while being well-versed in exotic species from online videos; we may enjoy cooking videos about exotic ingredients but remain unaware of unique vegetables like purple broccoli available at the local supermarket.
Emerging technologies like Multimodal Large Language Models (MLLMs) (Yin et al., 2023) and wearable devices, such as smart glasses, offer new possibilities for ”having new eyes” to discover hidden knowledge in our environment. By using cameras on smart glasses to sense the surroundings, MLLMs, with their vast knowledge base, can identify relevant information about entities in the environment and proactively deliver it via smart glasses, even when users are engaged in other tasks like commuting. However, this approach faces two challenges: (1) given the multitude of elements that can be explored in our surrounding environment, it remains unclear which types of knowledge users value most in everyday moments; (2) although proactive knowledge delivery may reduce the effort for manual searches (e.g., using Google or ChatGPT), it could also disrupt primary tasks if without careful design. These challenges may hinder the adoption of proactive wearable knowledge discovery assistants. As a result, existing work has largely focused either on user-initiated queries (e.g., users look at a target and ask for information) (Wang et al., 2024b; Lee et al., 2024; Google, 2024), or limited learning content, such as foreign vocabulary (Arakawa et al., 2022) or pre-registered material (e.g., AR in museums (Sommerauer and Müller, 2014)), limiting deeper engagement with the everyday environment.
@@
AiGet’s design is carefully tailored for everyday mobile scenarios, ensuring seamless integration alongside primary tasks, with two key aspects: 1) Information Content: To enhance perceived usefulness and reduce annoyance, AiGet prioritizes knowledge that balances novelty, personalization, usefulness, and unexpectedness (Kelly and Sharot, 2021; Sharot and Sunstein, 2020; Kotkov et al., 2023). 2) User Interface Design: To reduce interference and maintain environmental awareness, the system employs multimodal output: keywords and emojis for gist, images with bounding boxes for spatial reference, and audio for details, making knowledge absorption easier. To provide user control while minimizing conflicts during undesired learning moments (e.g., rushing through time-sensitive errands), AiGet uses a ring mouse and subtle gestures for interactions such as turn on/off, canceling displays, or initiating follow-up queries. This integrated design naturally embeds learning opportunities into daily activities while remaining unobtrusive and user-friendly.
We conducted in-lab simulations with 12 participants and real-world evaluations involving 18 participants (with 6 for continued usage for up to 7 days) in low-demand daily activities [^1] In our context, we focused on daily tasks with relatively low temporal or cognitive pressure, excluding tasks such as focused work to avoid overwhelming users., such as casual walking, shopping, and museum exploration. AiGet demonstrated effectiveness in uncovering overlooked knowledge in environments, providing information that expands users’ knowledge base while being useful, personalized, and surprising. As a result, it enhanced primary task enjoyment through relevant and timely information delivery, revived users’ attention and curiosity about surroundings, deepened their connection with their environment, and demonstrated potential for long-term usability across diverse daily contexts.
@@ 2. Related Work
This section provides (1) an overview of learning in daily life and the computing technologies that support it, (2) recent trends in using AI to enhance learning in everyday contexts, and (3) wearable solutions that facilitate learning during daily activities.
@@ 2.1. Computing Technologies for Learning in Daily Life
Despite these advances, traditional informal learning systems are often topic- or area-specific, such as foreign languages (Hautasaari et al., 2020), courses (Brahimi and Sarirete, 2015), or certain daily activities like cooking (Yanai et al., 2014), restricting their applicability to various everyday scenarios. We aim to extend this by utilizing context-aware devices (e.g., AR smart glasses) to support open-world knowledge discovery, addressing constraints such as time and topic limitations through AI-driven technologies.
2.2. Leveraging ML/AL to Enhance Daily Learning
@@
To enhance user experiences, LLM-based systems increasingly incorporate personalization through individual user profiles or group personas (Xu et al., 2024; Sun et al., 2024; Ha et al., 2024; Richardson et al., 2023). Personalized systems have been shown to increase engagement and motivation for informal learning (Draxler et al., 2022; Chen et al., 2015), yet many LLM/MLLM-based informal learning tools still lack personalization capabilities (with notable exceptions like the proposed G-VOILA framework (Wang et al., 2024b), yet without implementing this feature for real-world evaluation). While we envision integrating personalization with LLM-based systems could enhance informal knowledge generation, how to accommodate both long-term preferences and real-time context remains underexplored, especially for incidental learning scenarios where users lack explicit learning intentions. In this work, we investigate how personalized MLLM-based systems can leverage context-awareness to enhance the informal learning experience during daily activities.
@@ 3. Study Overview
Our research began with a formative study to understand user knowledge acquisition needs in daily life. Following this, we developed the proof-of-concept system, AiGet. We then conducted an in-lab evaluation to examine the AiGet system’s ability to generate desirable knowledge and a real-world study to test its feasibility and collect additional design insights. Note that all studies were approved by our university’s institutional review board (IRB), and participants were compensated $\approx$ 7.5 USD per hour, a standard rate for user studies in the local context.
@@ 4.1. Participants
We recruited twelve volunteers (P1-P12, 7 females, 5 males, Age: $M=24.9,\>SD=3.8$ years) from the university community, all with self-reported professional working fluency in English. To ensure the accuracy of our eye-tracking equipment, we selected participants with normal or corrected vision, excluding those wearing spectacles. Among the twelve participants, six reported using VQA apps or search engines (e.g., Google Lens) to query objects/entities in daily life at least twice per day, five used such tools with medium frequency (3-7 times per week), and one participant rarely used these technologies.
@@ 4.3. Study Design and Procedure
The study was conducted in three phases: 1) Pre-Study Interview: Participants were interviewed to understand their daily routines, informal learning habits, challenges, and desired learning goals. 2) Natural Behavior Recording: Participants followed a designated path (indoor to outdoor campus to a market with unfamiliar imported products) while their behaviors were recorded using the apparatus (Sec 4.2), without experimenter intervention. 3) Think-Aloud and Reflection: Participants retraced the path, thinking aloud about the knowledge they wanted and why. After completing the path, they reviewed their FPV and gaze recordings at 2x speed, annotating desired knowledge points (Janaka et al., 2024a). To understand if any knowledge was ignored yet desirable, a “Wizard of Oz” simulation using GPT-4V provided overlooked knowledge for familiar and unfamiliar entities (e.g., animals, plants, products) and focused and ignored items. Participants’ expectations for wearable AI interactions were also documented.
@@ 4.5.1. Four Common Barriers Limiting Users from Discovering Informal Knowledge in Daily Activities
Participants, primarily university students and staff, described weekday routines centered on commuting, work, and occasional sports, with weekends involving more personal activities like shopping and traveling. Within these routines, participants identified four main barriers to learning in their daily environments. (1) Time and Attention Limitation (Bruya, 2010; Baron, 2023): Aligning with informal learning literature (Marsick and Watkins, 2001; Livingstone, 2001b), all participants mentioned this as the most significant barrier. They are aware of missing out on new information because they focus on existing tasks and feel that the effort required to seek out new information outweighs the potential benefits, leading to ”Known Unknowns” [^3] Things we are aware of but don’t understand, from Rumsfeld Matrix (Saravanan, 2021). (2) Inattentional Blindness (Mack, 2003): Participants often failed to notice interesting elements in their environment due to cognitive constraints, such as being preoccupied with other tasks or work-related thoughts, leading to ”Unseen” knowledge. (3) Illusion of Explanatory Depth (Rozenblit and Keil, 2002): Participants overestimate their understanding or familiarity with surrounding entities, leading to ”Unknown Unknowns” [^4] Things we are neither aware of nor understand unless pointed out by others (e.g., “I never knew the commonly seen tree is the national tree here and people used it for cutting boards. (P2)” (4) Uncertainty Avoidance (Money and Crotts, 2003): Barriers like language gaps lead users to avoid exploring unfamiliar environments or subjects (e.g., “unfamiliar imported foods in stores”), also resulting in ”Known Unknowns” [^5] Things we are aware of but don’t understand. Although these barriers hinder knowledge discovery in daily life, P6 noted that they do not signify a loss of curiosity. Instead, they tend to make quick assumptions about unfamiliar things (e.g., unfamiliar fruit taste) to save time, even if those assumptions are sometimes incorrect.
Meanwhile, participants have shifted their primary informal knowledge acquisition to passive consumption through social media platforms like Twitter and YouTube, aligning with prior research (Oeldorf-Hirsch, 2018; Haythornthwaite, 2022). This behavior contrasts with their past, particularly childhood, when they actively questioned their surroundings. Now, they only seek in-situ knowledge under specific conditions: when highly interested and with ample free time (e.g., “I used plant recognition apps to search for very unique flowers I encountered (P9)”), or when the information has high practical value (e.g., “I will search for tips before traveling or going to exhibitions (P2)”).
@@ 4.5.2. Desired Types of Knowledge for Addressing Informal Learning Barriers
For the remaining three categories, we found their preferences are linked to the entity’s familiarity level, echoing the need to address Illusion of Explanatory Depth and Uncertainty Avoidance (Sec 4.5.1).
@@ 4.5.3. Summarized Design Requirements for Wearable AI Assistant to Help Users Learn from Environment
Participants appreciated discovering information from their environment but emphasized the need to prioritize the most valuable content, avoiding information that is trivial. To address this, we propose a multi-factor weighting system to prioritize knowledge selection. First, Novelty was considered crucial, with most participants (10 out of 12) agreeing that the knowledge must be new to them, either by introducing entirely new perspectives or by adding unknown details to common knowledge. For example, instead of simply stating that ”palm trees provide habitats for various species,” the assistant could highlight that ”palm trees provide habitats for species like the Palm Weevil, which lays its eggs in the trunk.” Additionally, participants identified three factors as beneficial, though not always necessary: (1) Alignment with Personal Interests and Values, such as providing bargain-hunting tips for participants who value “saving”; (2) Usefulness for ongoing or future tasks, like helping health-oriented users compare sugar levels when buying drinks; and (3) Unexpected Perspectives that offer surprising insights or correct misconceptions, such as ”zero-sugar drinks doesn’t mean it contains zero calories.”
@@
Given the use of the wearable assistant in scenarios where users have other primary tasks and consume information on the move, participants highlighted two main points to minimize distraction. First, regarding knowledge delivery, participants emphasized the need to minimize the cognitive load during receptive moments. As P7 noted, “the content should be concise and easy to digest, like a podcast.” They also emphasized that the assistant should provide only 1-2 pieces of information at a time to prevent overwhelming users, especially during tasks requiring focus, unless they explicitly requested more. Second, concerning system control, participants emphasized the importance of accommodating moments when users are unwilling to receive knowledge. This includes allowing users to turn the assistant off as needed for extended periods (e.g., while rushing to a destination or preoccupied with work or study) or easily interrupt single information delivery if the content is unsatisfactory. These considerations aim to ensure that the assistant enhances user experience rather than hinders the user’s primary activities in mobile settings.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 01, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 01, 2025, Yokohama, Japandoi: 10.1145/3706598.3713953isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Human-centered computing Empirical studies in interaction design
Figure 1. A day with AiGet, a wearable knowledge discovery assistant equipped with AR smart glasses and a ring mouse. (1) While casually walking, AiGet actively analyzes the user’s gaze patterns and identifies “interesting” environmental entities. (2) AiGet recognizes the spider lily that the user briefly glanced at as a valuable opportunity for informal learning, considering her interest in plants in the User Profile and lack of knowledge about the flower. (3) Using contextual data, the system generates multiple knowledge candidates and selects the most valuable one about the spider lily for the user. (4) AiGet delivers multimodal feedback, including audio, text keywords with emojis, and an image with a bounding box for the flower to help the user locate the flower and gain information with minimal interference to primary tasks. (5) Surprised by the knowledge delivery about the overlooked flower, the user uses the ring mouse to ask follow-up questions. (6) Later, AiGet continuously helps her uncover design details behind the campus architecture and compare two unfamiliar snacks, aligning with her interest in healthy diets, etc. (7) When the user meets a friend who owns a cat, she shares what she learned from AiGet—that spider lilies are toxic to pets—enriching their social interaction.
@@ 1. Introduction
Our daily surroundings are filled with rich knowledge, from building designs to local flora and fauna, offering countless opportunities for discovery and informal learning from everyday moments (Kaplan, 1989; Bekerman et al., 2006; Ardoin and Heimlich, 2021). Such informal learning not only allows us to gather useful knowledge about our surroundings but also fosters critical thinking and diverse perspectives, helping us adapt to new situations and inspire innovation (Marsick and Watkins, 2001; Cross, 2007; Eraut *, 2004). Yet, modern adult life often disconnects us from these rich opportunities, making us “blind” to such discoveries. Pressed by daily demands, adults focus on primary tasks with higher priority (e.g., rushing for work and studies), bypassing opportunities for curiosity and exploration. We also tend to prioritize immediate rewards, avoiding the perceived effort of learning (Liquin and Gopnik, 2022). Moreover, the rise of smartphones and social media, which offer effortless access to knowledge, has further diverted our attention from the physical world. As a result, we may overlook rare migratory birds in our backyard while being well-versed in exotic species from online videos; we may enjoy cooking videos about exotic ingredients but remain unaware of unique vegetables like purple broccoli available at the local supermarket.
Emerging technologies like Multimodal Large Language Models (MLLMs) (Yin et al., 2023) and wearable devices, such as smart glasses, offer new possibilities for ”having new eyes” to discover hidden knowledge in our environment. By using cameras on smart glasses to sense the surroundings, MLLMs, with their vast knowledge base, can identify relevant information about entities in the environment and proactively deliver it via smart glasses, even when users are engaged in other tasks like commuting. However, this approach faces two challenges: (1) given the multitude of elements that can be explored in our surrounding environment, it remains unclear which types of knowledge users value most in everyday moments; (2) although proactive knowledge delivery may reduce the effort for manual searches (e.g., using Google or ChatGPT), it could also disrupt primary tasks if not carefully designed. These challenges may hinder the adoption of proactive wearable knowledge discovery assistants. As a result, existing work has largely focused either on user-initiated queries (e.g., users look at a target and ask for information) (Wang et al., 2024b; Lee et al., 2024; Google, 2024), or limited learning content, such as foreign vocabulary (Arakawa et al., 2022) or pre-registered material (e.g., AR in museums (Sommerauer and Müller, 2014)), limiting deeper engagement with the everyday environment.
@@
AiGet’s design is carefully tailored for everyday mobile scenarios, ensuring seamless integration alongside primary tasks, with two key aspects: 1) Information Content: To enhance perceived usefulness and reduce annoyance, AiGet prioritizes knowledge that balances novelty, personalization, usefulness, and unexpectedness (Kelly and Sharot, 2021; Sharot and Sunstein, 2020; Kotkov et al., 2023). 2) User Interface Design: To reduce interference and maintain environmental awareness, the system employs multimodal output: keywords and emojis for gist, images with bounding boxes for spatial reference, and audio for details, making knowledge absorption easier. To provide user control while minimizing conflicts during undesired learning moments (e.g., rushing through time-sensitive errands), AiGet uses a ring mouse and subtle gestures for interactions such as turning the system on/off, canceling displays, or initiating follow-up queries. This integrated design naturally embeds learning opportunities into daily activities while remaining unobtrusive and user-friendly.
We conducted in-lab simulations with 12 participants and real-world evaluations involving 18 participants (with 6 for continued usage for up to 7 days) in low-demand daily activities [^1] In our context, we focused on daily tasks with relatively low temporal or cognitive pressure, excluding tasks such as focused work to avoid overwhelming users., such as casual walking, shopping, and museum exploration. AiGet demonstrated effectiveness in uncovering overlooked knowledge in environments, providing information that expands users’ knowledge bases while being useful, personalized, and surprising. As a result, it enhanced primary task enjoyment through relevant and timely information delivery, revived users’ attention and curiosity about surroundings, deepened their connection with their environment, and demonstrated potential for long-term usability across diverse daily contexts.
@@ 2. Related Work
This section provides: (1) an overview of learning in daily life and the computing technologies that support it, (2) recent trends in using AI to enhance learning in everyday contexts, and (3) wearable solutions that facilitate learning during daily activities.
@@ 2.1. Computing Technologies for Learning in Daily Life
Despite these advances, traditional informal learning systems are often topic- or domain-specific, such as foreign languages (Hautasaari et al., 2020), courses (Brahimi and Sarirete, 2015), or certain daily activities like cooking (Yanai et al., 2014), restricting their applicability to diverse everyday scenarios. We aim to extend this by utilizing context-aware devices (e.g., AR smart glasses) to support open-world knowledge discovery, addressing constraints such as time and topic limitations through AI-driven technologies.
2.2. Leveraging Machine Learning and AI to Enhance Daily Learning
@@
To enhance user experiences, LLM-based systems increasingly incorporate personalization through individual user profiles or group personas (Xu et al., 2024; Sun et al., 2024; Ha et al., 2024; Richardson et al., 2023). Personalized systems have been shown to increase engagement and motivation for informal learning (Draxler et al., 2022; Chen et al., 2015), yet many LLM/MLLM-based informal learning tools still lack personalization capabilities (with notable exceptions like the proposed G-VOILA framework (Wang et al., 2024b), yet without implementing this feature for real-world evaluation). While we envision integrating personalization with LLM-based systems could enhance informal knowledge generation, how to accommodate both long-term preferences and real-time context remains underexplored, especially for incidental learning scenarios where users lack explicit learning intentions. In this work, we investigate how personalized MLLM-based systems can leverage context awareness to enhance the informal learning experience during daily activities.
@@ 3. Study Overview
Our research began with a formative study to understand users’ knowledge acquisition needs in daily life. Subsequently, we developed the proof-of-concept system, AiGet. We then conducted an in-lab evaluation to assess AiGet’s ability to generate desirable knowledge and a real-world study to evaluate its feasibility and collect additional design insights. All studies were approved by our university’s institutional review board (IRB), and participants were compensated $\approx$ 7.5 USD per hour, a standard rate for user studies in the local context.
@@ 4.1. Participants
We recruited twelve volunteers (P1-P12, 7 females, 5 males, Age: $M=24.9,\>SD=3.8$ years) from the university community, all self-reporting professional working fluency in English. To ensure the accuracy of our eye-tracking equipment, we selected participants with normal or corrected vision, excluding those wearing spectacles. Among the twelve participants, six reported using VQA apps or search engines (e.g., Google Lens) to query objects/entities in daily life at least twice per day, five used such tools with medium frequency (3-7 times per week), and one participant rarely used these technologies.
@@ 4.3. Study Design and Procedure
The study was conducted in three phases: 1) Pre-Study Interview: Participants were interviewed to understand their daily routines, informal learning habits, challenges, and desired learning goals. 2) Natural Behavior Recording: Participants followed a designated path (indoor to outdoor campus to a market with unfamiliar imported products) while their behaviors were recorded using the apparatus (Sec 4.2), without experimenter intervention. 3) Think-Aloud and Reflection: Participants retraced the path, thinking aloud about the knowledge they wanted and why. After completing the path, they reviewed their FPV and gaze recordings at 2x speed, annotating desired knowledge points (Janaka et al., 2024a). To understand if any knowledge was ignored yet desirable, a “Wizard of Oz” simulation using GPT-4V provided overlooked knowledge for familiar and unfamiliar entities (e.g., animals, plants, products) and both attended and ignored items. Participants’ expectations for wearable AI interactions were also documented.
@@ 4.5.1. Four Common Barriers Limiting Users from Discovering Informal Knowledge in Daily Activities
Participants, primarily university students and staff, described weekday routines centered on commuting, work, and occasional sports, with weekends involving more personal activities like shopping and traveling. Within these routines, participants identified four main barriers to learning in their daily environments. (1) Time and Attention Limitation (Bruya, 2010; Baron, 2023): Aligning with informal learning literature (Marsick and Watkins, 2001; Livingstone, 2001b), all participants mentioned this as the most significant barrier. They are aware of missing out on new information because they focus on existing tasks and feel that the effort required to seek out new information outweighs the potential benefits, leading to ”Known Unknowns” [^3] Things we are aware of but don’t understand, from Rumsfeld Matrix (Saravanan, 2021). (2) Inattentional Blindness (Mack, 2003): Participants often failed to notice interesting elements in their environment due to cognitive constraints, such as being preoccupied with other tasks or work-related thoughts, leading to ”Unseen” knowledge. (3) Illusion of Explanatory Depth (Rozenblit and Keil, 2002): Participants overestimate their understanding or familiarity with surrounding entities, leading to ”Unknown Unknowns” [^4] Things we are neither aware of nor understand unless pointed out by others (e.g., “I never knew the commonly seen tree is the national tree here and people used it for cutting boards. (P2)” (4) Uncertainty Avoidance (Money and Crotts, 2003): Barriers like language gaps lead users to avoid exploring unfamiliar environments or subjects (e.g., “unfamiliar imported food in stores”), also resulting in ”Known Unknowns” [^5] Things we are aware of but don’t understand. Although these barriers hinder knowledge discovery in daily life, P6 noted that they do not signify a loss of curiosity. Instead, they tend to make quick assumptions about unfamiliar things (e.g., unfamiliar fruit taste) to save time, even if those assumptions are sometimes incorrect.
Meanwhile, participants have shifted their primary informal knowledge acquisition to passive consumption through social media platforms like Twitter and YouTube, aligning with prior research (Oeldorf-Hirsch, 2018; Haythornthwaite, 2022). This behavior contrasts with their past, particularly in childhood, when they actively questioned their surroundings. Now, they only seek in-situ knowledge under specific conditions: when highly interested and with ample free time (e.g., “I used plant recognition apps to search for very unique flowers I encountered (P9)”), or when the information has high practical value (e.g., “I will search for tips before traveling or going to exhibitions (P2)”).
@@ 4.5.2. Desired Types of Knowledge for Addressing Informal Learning Barriers
For the remaining three categories, we found that their preferences are linked to the entity’s familiarity level, echoing the need to address Illusion of Explanatory Depth and Uncertainty Avoidance (Sec 4.5.1).
@@ 4.5.3. Summarized Design Requirements for Wearable AI Assistant to Help Users Learn from Environment
Participants appreciated discovering information from their environment but emphasized the need to prioritize the most valuable content, avoiding information that is trivial. To address this, we propose a multi-factor weighting system to prioritize knowledge selection. First, Novelty was considered crucial, with most participants (10 out of 12) agreeing that the knowledge must be new to them, either by introducing entirely new perspectives or by adding unknown details to common knowledge. For example, instead of simply stating that ”palm trees provide habitats for various species,” the assistant could highlight that ”palm trees provide habitats for species like the Palm Weevil, which lays its eggs in the trunk.” Additionally, participants identified three factors as beneficial, though not always necessary: (1) Alignment with Personal Interests and Values, such as providing bargain-hunting tips for participants who value “saving”; (2) Usefulness for ongoing or future tasks, like helping health-oriented users compare sugar levels when buying drinks; and (3) Unexpected Perspectives that offer surprising insights or correct misconceptions, such as ”zero-sugar drinks do not mean they contain zero calories.”
@@
Given the use of the wearable assistant in scenarios where users have other primary tasks and consume information on the move, participants highlighted two main points to minimize distraction. First, regarding knowledge delivery, participants emphasized the need to minimize the cognitive load during receptive moments for learning. As P7 noted, “the content should be concise and easy to digest, like a podcast.” They also emphasized that the assistant should provide only 1-2 pieces of information at a time to prevent overwhelming users, especially during tasks requiring focus, unless they explicitly requested more. Second, concerning system control, participants emphasized the importance of accommodating moments when users are unwilling to receive knowledge. This includes allowing users to turn the assistant off as needed for extended periods (e.g., while rushing to a destination or preoccupied with work or study) or easily interrupt single information delivery if the content is unsatisfactory. These considerations aim to ensure that the assistant enhances the user experience rather than hinders the user’s primary activities in mobile settings.